# thanks: Contact author:

### v1

@@ II.2 Learning Attention Scores with Quantum States
In the attention matrix $\mathbf{A}$, each element $a_{i,j}$ is the scaled dot product of the $i$-th query vector $\mathbf{q}_{i}$ and the $j$-th key vector $\mathbf{k}_{j}$, followed by a softmax operation (see Equation 5). In this work, we use quantum circuits to compute individual attention scores. We learn representations of the embedding vectors $\mathbf{z}_{i}$, query vectors $\mathbf{q}_{i}$, and key vectors $\mathbf{k}_{i}$ as quantum states, denoted |$\ket{q_{i}}$ and |$\ket{k_{j}}$, and determine their inner product $\braket{q_{i}}{k_{j}}$. The value matrix $\mathbf{V}$ and subsequent operations, however, remain classical. Figure 2 illustrates our hybrid quantum-classical self-attention framework. The next subsection details the quantum circuit used to compute $\mathbf{A}$’s matrix elements.
@@ II.2.1 Quantum Encoding of Token and Positional Information
Figure 2: Quantum self-attention layer combining $\mathbf{QK}^{\top}$ calculated with quantum circuits and a classically computed value matrix $\mathbf{V}$ (Equation 3). Quantum token embeddings $\mathbf{\theta_{e}}$, positional angles $\mathbf{\theta_{p}}$, and learnable parameters $\mathbf{\theta_{q}}$, $\mathbf{\theta_{k}}$ are used in the unitary $U$ as the circuit that evolves the quantum states depicted in Equations 712. Each quantum circuit produces an attention score, thus there are $\frac{n^{2}+n}{2}$ instances of $U$ with their respective angles to ensure a fully populated masked attention matrix. The expectation value of the Pauli-X observable on the ancilla qubit (equivalent to a Hadamard transform and measurement in the computational basis as shown in Equations 13 and 14) is obtained and represents the dot product between query and key vectors. The original transformer implementation [1] scales attention scores by $\frac{1}{\sqrt{d_{k}}}$ to maintain a variance of 1. Since the dot products herein are obtained from the expectations of the quantum subsystems, they are bound on the closed interval $[-1,1]$. To maintain a variance of 1, they must be scaled by $\sqrt{d_{k}}$. The scaled values are stored in the masked attention matrix, softmax is applied to the rows, and the resulting matrix is multiplied with $\mathbf{V}$.
@@ II.2.2 Learning Query and Key States
Analogous to the preparation of |$\ket{z_{i}}$ in Equation 7, |$\ket{z_{j}}$ is prepared with the difference being all PQCs are controlled by the ancilla qubit. This controlled preparation ensures the modified Hadamard test yields a real-valued dot product $\braket{q_{i}}{k_{j}}$ between query and key states. After preparing |$\ket{z_{j}}$, a controlled PQC $CU_{k}$ transforms it into |$\ket{k_{j}}$, applied only when the ancilla is in |$\ket{1}$, resulting in the state:
@@
$\bra{\Psi_{7}}\left(Z\otimes I^{\otimes\log{d}}\right)\ket{\Psi_{7}}=\text{Re}\braket{q_{i}}{k_{j}},$ (14)
@@
Figure 4: Quantum circuit used to create query and key states from a given quantum token and positional encoding to produce an attention score when measured. $U_{e_{i}}/U_{e_{j}}$ and $U_{p_{i}}/U_{p_{j}}$ are the unitaries that create the token and positional encoding of the $i$/$j$-th token into the quantum state. $U_{q}$ and $U_{k}$ are the unitaries to learn query and key representations of the quantum states containing token and positional information. The mathematical description of states $\Psi_{1}$ to $\Psi_{7}$ are found in Equations 7 to 13. The expectation value on the ancilla qubit yields the desired query-key dot product $\text{Re}\braket{q_{i}}{k_{j}}$.
@@ II.3.2 Simultaneous Perturbation Stochastic Approximation
where $\mathbf{\Delta}$ is an $m$-dimensional vector with elements randomly chosen as ±$\pm 1$, and the perturbation step $\epsilon$ is set to $0.01$ in this work.
@@ II.4.1 Computational Complexity
The time complexity of the classical calculation of the attention matrix A in Equation 5 is $\mathcal{O}(n^{2}d)$ stemming from the multiplication of Q and $\textbf{K}^{\top}$ which are matrices of dimensions $n\times d$ and $d\times n$, respectively. Under the assumption of efficient preparation of quantum states, this modified-Hadamard test approach produces each inner product $\braket{q_{i}}{k_{j}}$ in $\mathcal{O}(1)$ time compared to the classical $\mathcal{O}(d)$, leading to a complexity of $\mathcal{O}(n^{2})$ for the population of A. In practice, we prepare states with $\mathcal{O}(\log{d})$ depth, bringing the overall practical complexity to $\mathcal{O}(n^{2}\log{d})$. It is important to note that preparing quantum states in $\mathcal{O}(\log{d})$ time may result in highly structured states. Specifically, low depth circuits are likely to produce quantum states with sparse amplitudes or pattern-like structures, which could make direct comparisons to classical algorithms that are designed for dense and unstructured data less equitable.
@@ II.4.2 Query Complexity
The measurement of the ancilla qubit produces a Bernoulli random variable where the probability of measuring 1 is given by $p_{0}=\frac{1+\text{Re}\braket{q_{i}}{k_{j}}}{2}$. Thus, the Chernoff–Hoeffding theorem demonstrates $\text{Re}\braket{q_{i}}{k_{j}}$ can be found in $\mathcal{O}\left(\frac{1}{\epsilon^{2}}\right)$ query complexity with additive error $\epsilon$. However, since $\text{Re}\braket{q_{i}}{k_{j}}=2p_{0}-1$ is obtained by measuring the ancilla qubit in the |$\ket{0}$ state, quantum amplitude amplification can be performed to improve query complexity to $\mathcal{O}\left(\frac{1}{\epsilon}\right)$. To illustrate this, the final state before measurement as described in Equation 13 can be written equivalently:
@@ III.1 Dataset & Features
We employ the QM9 dataset [27], a benchmark of 133,885 small organic molecules represented as SMILES strings [28], as the basis for our study. The SMILES are canonicalized with RDKit [29] and duplicates were removed leaving 133,798 remaining datapoints. After preprocessing, the dataset was split into training and validation sets at a 20:1 ratio. From the dataset, we extracted nine key physicochemical features using RDKit: molecular weight (MW), number of hydrogen bond acceptors (HBA), number of hydrogen bond donors (HBD), number of rotatable bonds (nRot), number of rings (nRing), number of heteroatoms (nHet), topological polar surface area (TPSA), logP (partition coefficient), and the number of stereocenters (Stereo). To incorporate these descriptors as additional embeddings into our quantum-classical hybrid model, we transform them via a classical linear layer from 9 dimensions to $\dim{\theta_{e}}=\dim{\theta_{p}}$. We then scale the batch linearly between between 00 and $\pi$ to produce $\theta_{c}$ for quantum circuit encoding. The SMILES strings are tokenized by breaking them into meaningful substructures, such as atoms, rings, branches, and bond types, and converting them into a sequence of discrete tokens that can be mapped to a high-dimensional embedding vector. The QM9 SMILES strings consist of 30 unique tokens, along with padding, start-of-sequence, and end-of-sequence tokens, resulting in a total vocabulary size of 33.
@@ III.2 Benchmarks & Trainings
We compared the quantum-classical model’s performance to that of fully classical models with equivalent architectures. All training setups employed one decoder layer and one attention head. The quantum model computed attention scores using 6 active qubits, producing a Hilbert space of dimension $642^{6}=$ 64, and thus we set the classical token and positional embedding vectors to 64 dimensions. We also trained a classical model with equal parameter counts, denoted Classical–eq, for further comparison. Since each PQC ansatz uses a single layer of $R_{y}$ gates, the number of learnable parameters per register matches the qubit count—3 for token and positional registers in sequence-only training ($\frac{\log 64}{2}=3$, Equation 6), and 2 for token, positional, and physicochemical registers in condition-based training ($\frac{\log 64}{3}=2$, Equation 15). To match the parameter count between the Quantum and Classical–eq models, the weight matrices $\mathbf{W}^{Q}$ and $\mathbf{W}^{K}$ (Equation 3) have shapes $3\times 2$ and $2\times 3$ for sequence-only and condition-based setups, respectively, yielding 6 parameters each for query and key transformations. All models in this work use 64-dimensional embeddings for the value matrix $\mathbf{W}^{V}$. To summarize, the total number of parameters for the hybrid quantum-classical model (Quantum), fully classical model with an equal number of parameters (Classical–eq) model were 47,704 and 48,307 for the sequence and condition-based models, respectively. The fully classical model with an equivalent architecture to the Quantum model but with traditionally sized query-key weight matrices $\mathbf{W}^{Q}$ and $\mathbf{W}^{K}$ of shape $64\times 64$ (denoted as Classical) has 55,713 and 56,535 parameters for the sequence and condition-based model, respectively. To fairly evaluate the performance of these models, we initialized shared parameters across all models with identical random values.
@@
We implemented the machine learning components using PyTorch [30]. All models underwent training for 20 epochs with the AdamW optimizer [31], set to a learning rate of 0.005 and weight decay of 0.1. We applied gradient clipping with a maximum norm of 1.0 per layer to stabilize gradients and used cross-entropy loss as the objective function. To support a batch size of 256, we conducted quantum circuit simulations with CUDA-Q [32], an open-source QPU-agnostic platform designed for accelerated quantum-classical supercomputing. All quantum simulations were performed using the state-vector simulator available in CUDA-Q. Training times for a single epoch on a CPU versus a single GPU are shown in Table 1, with a single GPU achieving a speedup of 1.34 over the CPU. Distributing simulations across four GPUs further yielded a speedup of 3.84 relative to one GPU. Consequently, we utilized four NVIDIA A100 GPUs on NERSC’s Perlmutter supercomputer to accelerate training with this large batch size.
Table 1: The training time for a single epoch on the QM9 dataset with the condition-based Quantum model using a batch size of 256. Runtimes for the full training are extrapolated from the average runtime over 4 batch updates. The GPUs and CPU used here are NVIDIA A100s and an AMD EPYC 7763, respectively.
Hardware Epoch Time (hrs)
@@ IV.3 Out-of-Distribution Modeling Performance
For targets $2\sigma$ above the mean, the quantum model generated molecules closest to the targets for four properties: nRing, nHet, LogP, and Stereo. The Classical–eq model matched three properties (MW, HBD, TPSA), while the Classical model outperformed others on two (HBA, nRot). Below the mean by $2\sigma$, the quantum model excelled at HBD, nRot, TPSA, and LogP; the Classical–eq model at MW and nHet; and the Classical model at HBA and Stereo. Notably, all models generated molecules with zero rings equally well. To assess whether skewed distributions disproportionately affect any model, we repeated the experiment using median ±$\pm 1.5$ IQR targets, where IQR is the interquartile range. Results in Table A2 confirm further that all models exhibit comparable performance in generating molecules beyond the training distribution.
@@ Appendix A Property Choices for Inference
For all models (Quantum, Classical–eq, Classical), the ranking of property value selection methods that maximize valid and unique molecules ($V\times U$) is: Mean $>$ Mode $>$ Median, as shown in Table A1. Interestingly, this statistical choice affects novelty, with mode-based inference increasing the fraction of novel compounds to over 95%. We tested the models’ ability to generate molecules beyond the training distribution using $\text{median}\pm 1.5$ IQR targets, as presented in Table A2, to assess whether distributional skew disproportionately impacts any model. For the upper range, the Quantum model performed best for nRing; the Classical–eq model for MW, nHet, LogP, and Stereo; and the Classical model for HBA, HBD, nRot, and TPSA. For the lower range, the Quantum model performed best for HBA, nHet, and LogP; the Classical–eq model for TPSA; and the Classical model for MW, HBD, nRot, and Stereo. These results–given the margins observed–suggest no model outperforms or underperforms in generating target molecules when using $\mu\pm 2\sigma$ versus $\text{median}\pm 1.5$ IQR approaches.

### v2

@@ II.2 Learning Attention Scores with Quantum States
In the attention matrix $\mathbf{A}$, each element $a_{i,j}$ is the scaled dot product of the $i$-th query vector $\mathbf{q}_{i}$ and the $j$-th key vector $\mathbf{k}_{j}$, followed by a softmax operation (see Equation 5). In this work, we use quantum circuits to compute individual attention scores. We learn representations of the embedding vectors $\mathbf{z}_{i}$, query vectors $\mathbf{q}_{i}$, and key vectors $\mathbf{k}_{i}$ as quantum states, denoted |$\ket{q_{i}}$ and |$\ket{k_{j}}$, and determine their inner product $\braket{q_{i}|k_{j}}$. The value matrix $\mathbf{V}$ and subsequent operations, however, remain classical. Figure 2 illustrates our hybrid quantum-classical self-attention framework. The next subsection details the quantum circuit used to compute $\mathbf{A}$’s matrix elements.
@@ II.2.1 Quantum Encoding of Token and Positional Information
Figure 2: Quantum self-attention layer combining $\mathbf{QK}^{\top}$ calculated with quantum circuits and a classically computed value matrix $\mathbf{V}$ (Equation 3). Quantum token embeddings $\mathbf{\theta_{e}}$, positional angles $\mathbf{\theta_{p}}$, and learnable parameters $\mathbf{\theta_{q}}$, $\mathbf{\theta_{k}}$ are used in the unitary $U$ as the circuit that evolves the quantum states depicted in Equations 712. Each quantum circuit produces an attention score, thus there are $\frac{n^{2}+n}{2}$ instances of $U$ with their respective angles to ensure a fully populated masked attention matrix. The expectation value of the Pauli-X observable on the ancilla qubit (equivalent to a Hadamard transform and measurement in the computational basis as shown in Equations 13 and 14) is obtained and represents the dot product between query and key vectors. The original transformer implementation [1] scales attention scores by $\frac{1}{\sqrt{d_{k}}}$ to maintain a variance of 1. Since the dot products herein are obtained from the expectations of the quantum subsystems, they are bound on the closed interval [1,1][-1,1][ - 1 , 1 ]. To maintain a variance of 1, they must be scaled by $\sqrt{d_{k}}$. The scaled values are stored in the masked attention matrix, softmax is applied to the rows, and the resulting matrix is multiplied with $\mathbf{V}$.
@@ II.2.2 Learning Query and Key States
Analogous to the preparation of |$\ket{z_{i}}$ in Equation 7, |$\ket{z_{j}}$ is prepared with the difference being all PQCs are controlled by the ancilla qubit. This controlled preparation ensures the modified Hadamard test yields a real-valued dot product $\braket{q_{i}|k_{j}}$ between query and key states. After preparing |$\ket{z_{j}}$, a controlled PQC $CU_{k}$ transforms it into |$\ket{k_{j}}$, applied only when the ancilla is in |$\ket{1}$, resulting in the state:
@@
$\bra{\Psi_{7}}\left(Z\otimes I^{\otimes\log{d}}\right)\ket{\Psi_{7}}=\text{Re}\braket{q_{i}|k_{j}},$ (14)
@@
Figure 4: Quantum circuit used to create query and key states from a given quantum token and positional encoding to produce an attention score when measured. $U_{e_{i}}/U_{e_{j}}$ and $U_{p_{i}}/U_{p_{j}}$ are the unitaries that create the token and positional encoding of the $i$/$j$-th token into the quantum state. $U_{q}$ and $U_{k}$ are the unitaries to learn query and key representations of the quantum states containing token and positional information. The mathematical description of states $\Psi_{1}$ to $\Psi_{7}$ are found in Equations 7 to 13. The expectation value on the ancilla qubit yields the desired query-key dot product $\text{Re}\braket{q_{i}|k_{j}}$.
@@ II.3.2 Simultaneous Perturbation Stochastic Approximation
where $\mathbf{\Delta}$ is an $m$-dimensional vector with elements randomly chosen as ±1$\pm$ 1± 1, and the perturbation step $\epsilon$ is set to 0.010.010.01 in this work.
@@ II.4.1 Computational Complexity
The time complexity of the classical calculation of the attention matrix A in Equation 5 is $\mathcal{O}(n^{2}d)$ stemming from the multiplication of Q and $\textbf{K}^{\top}$ which are matrices of dimensions $n\times d$ and $d\times n$, respectively. Under the assumption of efficient preparation of quantum states, this modified-Hadamard test approach produces each inner product $\braket{q_{i}|k_{j}}$ in $\mathcal{O}(1)$ time compared to the classical $\mathcal{O}(d)$, leading to a complexity of $\mathcal{O}(n^{2})$ for the population of A. In practice, we prepare states with $\mathcal{O}(\log{d})$ depth, bringing the overall practical complexity to $\mathcal{O}(n^{2}\log{d})$. It is important to note that preparing quantum states in $\mathcal{O}(\log{d})$ time may result in highly structured states. Specifically, low depth circuits are likely to produce quantum states with sparse amplitudes or pattern-like structures, which could make direct comparisons to classical algorithms that are designed for dense and unstructured data less equitable.
@@ II.4.2 Query Complexity
The measurement of the ancilla qubit produces a Bernoulli random variable where the probability of measuring 1 is given by $p_{0}=\frac{1+\text{Re}\braket{q_{i}|k_{j}}}{2}$. Thus, the Chernoff–Hoeffding theorem demonstrates $\text{Re}\braket{q_{i}|k_{j}}$ can be found in $\mathcal{O}\left(\frac{1}{\epsilon^{2}}\right)$ query complexity with additive error $\epsilon$. However, since $\text{Re}\braket{q_{i}|k_{j}}=2p_{0}-1$ is obtained by measuring the ancilla qubit in the |$\ket{0}$ state, quantum amplitude amplification can be performed to improve query complexity to $\mathcal{O}\left(\frac{1}{\epsilon}\right)$. To illustrate this, the final state before measurement as described in Equation 13 can be written equivalently:
@@ III.1 Dataset & Features
We employ the QM9 dataset [27], a benchmark of 133,885 small organic molecules represented as SMILES strings [28], as the basis for our study. The SMILES are canonicalized with RDKit [29] and duplicates were removed leaving 133,798 remaining datapoints. After preprocessing, the dataset was split into training and validation sets at a 20:1 ratio. From the dataset, we extracted nine key physicochemical features using RDKit: molecular weight (MW), number of hydrogen bond acceptors (HBA), number of hydrogen bond donors (HBD), number of rotatable bonds (nRot), number of rings (nRing), number of heteroatoms (nHet), topological polar surface area (TPSA), logP (partition coefficient), and the number of stereocenters (Stereo). To incorporate these descriptors as additional embeddings into our quantum-classical hybrid model, we transform them via a classical linear layer from 9 dimensions to $\dim{\theta_{e}}=\dim{\theta_{p}}$. We then scale the batch linearly between between 0 and $\pi$ to produce $\theta_{c}$ for quantum circuit encoding. The SMILES strings are tokenized by breaking them into meaningful substructures, such as atoms, rings, branches, and bond types, and converting them into a sequence of discrete tokens that can be mapped to a high-dimensional embedding vector. The QM9 SMILES strings consist of 30 unique tokens, along with padding, start-of-sequence, and end-of-sequence tokens, resulting in a total vocabulary size of 33.
@@ III.2 Benchmarks & Trainings
We compared the quantum-classical model’s performance to that of fully classical models with equivalent architectures. All training setups employed one decoder layer and one attention head. The quantum model computed attention scores using 6 active qubits, producing a Hilbert space of dimension $2^{6}=64$, and thus we set the classical token and positional embedding vectors to 64 dimensions. We also trained a classical model with equal parameter counts, denoted Classical–eq, for further comparison. Since each PQC ansatz uses a single layer of $R_{y}$ gates, the number of learnable parameters per register matches the qubit count—3 for token and positional registers in sequence-only training ($\frac{\log 64}{2}=3$, Equation 6), and 2 for token, positional, and physicochemical registers in condition-based training ($\frac{\log 64}{3}=2$, Equation 15). To match the parameter count between the Quantum and Classical–eq models, the weight matrices $\mathbf{W}^{Q}$ and $\mathbf{W}^{K}$ (Equation 3) have shapes 3×23$\times$ 23 × 2 and 2×32$\times$ 32 × 3 for sequence-only and condition-based setups, respectively, yielding 6 parameters each for query and key transformations. All models in this work use 64-dimensional embeddings for the value matrix $\mathbf{W}^{V}$. To summarize, the total number of parameters for the hybrid quantum-classical model (Quantum), fully classical model with an equal number of parameters (Classical–eq) model were 47,704 and 48,307 for the sequence and condition-based models, respectively. The fully classical model with an equivalent architecture to the Quantum model but with traditionally sized query-key weight matrices $\mathbf{W}^{Q}$ and $\mathbf{W}^{K}$ of shape $\times 64$ 64 (denoted as Classical) has 55,713 and 56,535 parameters for the sequence and condition-based model, respectively. To fairly evaluate the performance of these models, we initialized shared parameters across all models with identical random values.
@@
We implemented the machine learning components using PyTorch [30]. All models underwent training for 20 epochs with the AdamW optimizer [31], set to a learning rate of 0.005 and weight decay of 0.1. We applied gradient clipping with a maximum norm of 1.0 per layer to stabilize gradients and used cross-entropy loss as the objective function. To support a batch size of 256, we conducted quantum circuit simulations with CUDA-Q [32], an open-source QPU-agnostic platform designed for accelerated quantum-classical supercomputing. All quantum simulations were performed using the state-vector simulator available in CUDA-Q. Training times for a single epoch on a CPU versus a single GPU are shown in Table 1, with a single GPU achieving a speedup of 1.34x over the CPU for the simulation with 7 qubits. The speedup here is moderate, as the size of the embedding is small and hence it does not fully saturate the GPU. Moreover, while the CPU time reflects pure computation, the GPU time includes data transfer overhead between host and device. For a larger embedding of a 13-qubit system, the GPU achieves a 5.75x speedup in runtime compared to the CPU. This gap is expected to widen further as we consider larger model sizes.
Distributing simulations across four GPUs results in a 3.84x speedup for the 7-qubit system and a 3.96x speedup for the 13-qubit system, compared to using a single GPU. Consequently, we utilized four NVIDIA A100 GPUs on NERSC’s Perlmutter supercomputer to accelerate training with this large batch size.
Table 1: Training time per epoch for the condition-based quantum model on the QM9 dataset using a batch size of 256. Values are extrapolated from the average runtime over four batch updates. All runs used an AMD EPYC 7763 CPU (64 cores, 128 threads, default thread settings for CUDA-Q and Pytorch) and NVIDIA A100 GPUs. For the 7-qubit model, GPU speedup over CPU is modest because the quantum circuits are relatively small, leading to limited GPU utilization and increased sensitivity to data transfer overhead between host and device. As system size increases, circuit complexity grows, allowing the GPU to more effectively leverage its parallel architecture. For the 13-qubit model, this results in a 5.75x speedup over CPU performance, and an additional 3.96x speedup when scaling across four GPUs.
Qubits Hardware Epoch Time (hrs)
@@ IV.3 Out-of-Distribution Modeling Performance
For targets $2\sigma$ above the mean, the quantum model generated molecules closest to the targets for four properties: nRing, nHet, LogP, and Stereo. The Classical–eq model matched three properties (MW, HBD, TPSA), while the Classical model outperformed others on two (HBA, nRot). Below the mean by $2\sigma$, the quantum model excelled at HBD, nRot, TPSA, and LogP; the Classical–eq model at MW and nHet; and the Classical model at HBA and Stereo. Notably, all models generated molecules with zero rings equally well. To assess whether skewed distributions disproportionately affect any model, we repeated the experiment using median ±1.5$\pm$ 1.5± 1.5 IQR targets, where IQR is the interquartile range. Results in Table A2 confirm further that all models exhibit comparable performance in generating molecules beyond the training distribution.
@@ Appendix A Property Choices for Inference
[…]